# Diffusive Gibbs Sampling

### v2

@@ 4 Empirical Evaluation
We evaluate DiGS on three complex multi-modal sampling tasks across various domains: a mixture of 40 Gaussians, Bayesian neural network, and molecular dynamics [^1] The code of our experiments can be found in https://github.com/Wenlin-Chen/DiGS.. For DiGS, we employ MALA with the MH initialization strategy for denoising sampling We compare DiGS with three baselines: MALA, HMC and PT. In all experiments, the step sizes of MALA and HMC are tuned via trial-and-error so that the acceptance rates are close to $0.574$ and $0.65$, respectively. We choose not to compare with RDMC, since it is computationally intractable on these complex tasks as demonstrated in Section 3.4.
@@ 4.1 Mixture of 40 Gaussians
Figure 8 shows a visual comparison for $10^{4}$ samples generated by each method. We can see that MALA and HMC fail to explore the modes that are far away from the origin. PT covers all 40 modes but produces significantly less samples for the modes on the top-right and bottom-right corners. DiGS manages to cover all modes with the right amount of samples in each mode. Table 2 shows the the Maximum Mean Discrepancy (MMD) (Gretton et al., 2012) (computed with 5 kernels with bandwidths $\{2^{-2},2^{-1},2^{0},2^{1},2^{2}\}$) between the true samples and samples generated by each sampler and the Mean Absolute Error (MAE) between the true and estimated expectations of a quadratic function under the MoG-40 target. This demonstrates that our method significantly outperforms all baselines on this problem.
@@ 4.2 Bayesian Neural Networks
In this experiment, all methods are initialized at the same random sample from the prior and generate $10^{3}$ samples from the posterior $p(\theta|D_{\text{train}})$ for evaluation. MAP runs $5\times 10^{6}$ full-batch gradient descent steps. MALA runs 5,000 Langevin steps per sample. HMC runs 5,000 leapfrog steps per sample. PT consists of 5 chains with temperatures $\tau=\{1.0,5.62,31.62,177.83,1000.0\}$, where each chain is constructed by an HMC sampler with 1,000 leapfrog steps per sample. DiGS uses the VP schedule with $T=5$ noise levels, ranging from $\alpha_{T}=0.1$ to $\alpha_{1}=0.9$, each with 100 Gibbs sweeps and 10 MALA denoising sampling steps per Gibbs sweep. Table 3 reports the average predictive negative log-likelihood (NLL) for each method on the test data, which shows that DiGS significantly outperforms other baselines. We speculate that the performance gain comes from the fact that DiGS captures a broader range of modes.
@@
MALA $0.335\pm 0.014$ $5.0\times 10^{6}$
HMC $0.317\pm 0.012$ $5.0\times 10^{6}$
PT $0.261\pm 0.005$ $5.0\times 10^{6}$
DiGS $\mathbf{0.199\pm 0.002}$ $5.0\times 10^{6}$
@@ 4.3 Molecular Dynamics
In this experiment, each method is initialized at the minimum energy configuration as in Midgley et al. (2023) and generates $10^{6}$ samples of configurations for evaluation. MALA runs 1,000 Langevin steps per sample. HMC runs 1,000 leapfrog steps per sample. DiGS uses the VP schedule with $T=2$ noise levels ($\alpha_{2}=0.1$ and $\alpha_{1}=0.9$), each with 100 Gibbs sweeps and 5 MALA denoising sampling steps per Gibbs sweep. Note that PT is the gold-standard method in molecular dynamics (MD) simulation, which serves as the ground-truth. The PT samples are taken from Midgley et al. (2023), which is generated using 21 chains starting at temperature 300K and increasing the temperature by 50K for each subsequent chain, where each chain is constructed by an HMC sampler with 1,000 leapfrog steps per sample. We follow Midgley et al. (2023) and treat $10^{7}$ PT samples as the ground-truth. In addition, we consider a baseline of $10^{6}$ PT samples as a reference. The quality of the sampled configurations is assessed by the Ramachandran plot (Ramachandran et al., 1963), which can be used to analyze how the protein folds locally. A Ramachandran plot is a 2D histogram for the two dihedral angles $\phi,\psi$ in the bonds connecting an amino acid to the protein backbone as shown in Figure 9. Table 4 shows the KL divergences for the Ramachandran plots and the marginals of $\phi,\psi$ between the ground-truth samples and samples generated by each sampler. We can see that DiGS significantly outperforms MALA and HMC with the same number of energy evaluations. DiGS also outperforms the gold-standard MD simulation method PT with 23×23$\times$23 × less energy evaluations. Figure 10 shows a visual comparison between the ground-truth and DiGS Ramachandran plots, confirming that DiGS captures all modes with the correct weightings.
@@ 5 Conclusion
Diffusive Gibbs Sampler offers significant improvements in sampling multi-modal distributions, surpassing traditional methods in both efficiency and accuracy. Its applicability across various fields (e.g., Bayesian inference and molecular dynamics) showcases the potential to facilitate sampling complex distributions in numerous scientific applications.
Acknowledgements and Disclosure of Funding
We thank Andi Zhang for the useful discussions on RDMC, Ruqi Zhang for highlighting Entropy-MCMC, and Arnaud Doucet, Michael Hutchinson, and Sam Power for pointing out proximal MCMC. MZ and DB acknowledge funding from the Cisco Centre of Excellence. WC acknowledges funding via a Cambridge Trust Scholarship (supported by the Cambridge Trust) and a Cambridge University Engineering Department Studentship (under grant G105682 NMZR/089 supported by Huawei R&D UK). JMHL acknowledges support from a Turing AI Fellowship under grant EP/V023756/1.
Reproducibility Statement
The implementation of our proposed DiGS method and the code of our experiments can be found in https://github.com/Wenlin-Chen/DiGS.
@@
$\log p_{\beta}(x=0)=\beta\left(-\mu^{2}/2\sigma_{g}^{2}-\log\sigma_{g}\right)+C_{\beta}.$ (20)
@@
Appendix C Experimental Details
In all experiments, the step sizes of MALA and HMC are tuned via trial-and-error so that the acceptance rates are close to optimal values $0.574$ and $0.65$, respectively.
C.1 Comparison of Initialization Strategies for the Denoising Sampling Step
C.2 Comparison of Gaussian Convolution Hyperparameters
C.3 Multi-Level Noise
For the multi-level noise experiment with the VP schedule in Section 2.3, we define $\alpha_{t}=\alpha_{T}+(T-t)\Delta_{\alpha}$, where $\Delta_{\alpha}=(\alpha_{1}-\alpha_{T})/(T-1)$ and $\sigma_{t}=\sqrt{1-\alpha_{t}}$. We set $\alpha_{1}=0.9$ and $\alpha_{T}=0.1$, and experiment with $T\in\{2,3,4,5\}$ for the comparison shown in Figure 4(c). For evaluation, we employ the Maximum Mean Discrepancy (MMD) (Gretton et al., 2012) that utilizes 5 kernels with bandwidths $\{2^{-2},2^{-1},2^{0},2^{1},2^{2}\}$.
C.4 Comparison with Parallel Tempering
For comparison with parallel tempering on the “Mixture of Delta” problem in Figure 6 in Section 3.1, each sampler is initialized at the origin and generates 1,000 samples. DiGS employs a single noise level Gaussian convolution with parameters $\alpha=1$ and $\sigma=1$, complemented by 5,000 Gibbs sweeps. Each Gibbs sweep begins with MH (Metropolis-Hastings) initialization, followed by 10 steps of MALA with a step size of $1{\times}10^{-3}$. PT consists of 5 chains with temperatures $\tau{=}\{1.0,5.62,31.62,177.83,1000.0\}$, where each chain is constructed by an HMC sampler with 1,000 leapfrog steps per sample and a step size of $1{\times}10^{-2}$.
C.5 Comparison with RDMC
For the comparison with Reverse Diffusion Monte Carlo (RDMC) in Figure 7 in Section 3.4, we follow the original setting discussed in the paper (Huang et al., 2023) and use ULA for generating samples for the posterior distribution. The score function of the reverse distribution using $K=5$ to construct the score approximation
$\nabla_{x_{t}}\log p(x_{t})\approx\frac{\alpha}{K}\sum_{k=1}^{K}\left(x_{0}^{(t,k)}-x_{t}\right)/\sigma_{t}^{2},$ (22)
where $x_{0}^{(t,k)}\sim p(x_{0}|x_{t})$ denotes the samples drawn using ULA with $L_{ULA}=5$ steps and a step size of $1{\times}10^{-2}$. These samples are initialized through importance sampling (IS) with $S_{is}=100$ samples. The discretization step size $\eta$ is set to $\Gamma/T$, with the scaling factor $a_{t}=e^{-(\Gamma-t\cdot\eta)}$ and the standard deviation $\sigma_{t}=\sqrt{1-\alpha_{t}^{2}}$. Following the procedure in (Huang et al., 2023), we further apply ULA for additional $L_{ULA}$ steps after RDMC. The parameter $\Gamma$ is fixed at $0.1$ for our experiments, and we explore different values of $T\in\{1,2,3,4\}$, necessitating $TK(L_{ULA}+S_{is})+L_{ULA}$ energy evaluations to generate a single sample.
C.6 Mixture of 40 Gaussians
C.7 Bayesian Neural Networks
For the BNN experiment in Section 4.2, all methods are initialized at the same random sample from the prior and generate $10^{3}$ samples from the posterior $p(\theta|D_{\text{train}})$ for evaluation. MAP runs $5\times 10^{6}$ full-batch gradient descent steps with a step size of $3\times 10^{-2}$. MALA runs 5,000 Langevin steps per sample with a step size of $1\times 10^{-4}$. HMC runs 5,000 leapfrog steps per sample with a step size of $5\times 10^{-4}$. PT consists of 5 chains with temperatures $\tau=\{1.0,5.62,31.62,177.83,1000.0\}$, where each chain is constructed by an HMC sampler with 1,000 leapfrog steps per sample with a step size of $5\times 10^{-4}$. DiGS uses the VP schedule with $T=5$ noise levels, ranging from $\alpha_{T}=0.1$ to $\alpha_{1}=0.9$, each with 100 Gibbs sweeps and 10 MALA denoising sampling steps per Gibbs sweep with a step size of $1\times 10^{-4}$.
C.8 Molecular Dynamics
For the molecular dynamics experiment in Section 4.3, each method is initialized at the minimum energy configuration as in Midgley et al. (2023) and generates $10^{6}$ samples of configurations for evaluation. MALA runs 1,000 Langevin steps per sample with a step size of $1\times 10^{-4}$. HMC runs 1,000 leapfrog steps per sample with a step size of $1\times 10^{-3}$. DiGS uses the VP schedule with $T=2$ noise levels ($\alpha_{2}=0.1$ and $\alpha_{1}=0.9$), each with 100 Gibbs sweeps and 5 MALA denoising sampling steps per Gibbs sweep with a step size of $1\times 10^{-4}$. Note that PT is the gold-standard method in molecular dynamics (MD) simulation, which serves as the ground-truth. The PT samples are taken from Midgley et al. (2023), which is generated using 21 chains starting at temperature 300K and increasing the temperature by 50K for each subsequent chain, where each chain is constructed by an HMC sampler with 1,000 leapfrog steps per sample. We follow Midgley et al. (2023) and treat $10^{7}$ PT samples as the ground-truth. In addition, we consider a baseline of $10^{6}$ PT samples as a reference.
Appendix D Extra Experimental Results
Figure 11 shows the Ramachandran plots produced by all compared methods for the molecular dynamics experiment in Section 4.3.

### v5

@@ 4 Empirical Evaluation
We evaluate DiGS on three complex multi-modal sampling tasks across various domains [^1] The code of our experiments can be found in https://github.com/Wenlin-Chen/DiGS.: a mixture of 40 Gaussians, Bayesian neural network, and molecular dynamics. For DiGS, we employ MALA with the Metropolis-within-Gibbs scheme. We compare DiGS with three baselines: MALA, HMC and PT. In all experiments, the step sizes of MALA and HMC are tuned via trial-and-error so that the acceptance rates are close to $0.574$ (Roberts & Rosenthal, 1998) and $0.65$ (Neal et al., 2011), respectively. We choose not to compare with RDMC, since it is computationally intractable on these complex tasks as demonstrated in Section 3.4.
@@ 4.1 Mixture of 40 Gaussians
Figure 8 shows a visual comparison for $10^{4}$ samples generated by each method. We can see that MALA and HMC fail to explore the modes that are far away from the origin. PT covers all 40 modes but produces significantly less samples for the modes on the top-right and bottom-right corners. DiGS manages to cover all modes with the right amount of samples in each mode. Table 2 shows the Maximum Mean Discrepancy (MMD) (Gretton et al., 2012) (computed with 5 kernels with bandwidths $\{2^{-2},2^{-1},2^{0},2^{1},2^{2}\}$) between the true samples and samples generated by each sampler and the Mean Absolute Error (MAE) between the true and estimated expectations of a quadratic function under the MoG-40 target. This demonstrates that our method significantly outperforms all baselines on this problem.
@@ 4.2 Bayesian Neural Networks
All methods are initialized at the same random sample from the prior and generate $150$ samples from the posterior $p(\theta|D_{\text{train}})$ for evaluation. MAP (Maximum a Posteri) runs $7.5\times 10^{5}$ full-batch gradient descent steps. MALA runs 5,000 Langevin steps per sample. HMC runs 5,000 leapfrog steps per sample. PT consists of 5 chains with temperatures $\tau=\{1.0,5.62,31.62,177.83,1000.0\}$, where each chain is constructed by an HMC sampler with 1,000 leapfrog steps per sample. DiGS uses the VP schedule with $T=5$ noise levels, ranging from $\alpha_{T}=0.1$ to $\alpha_{1}=0.9$, each with 100 Gibbs sweeps and 10 MALA denoising sampling steps per Gibbs sweep. Table 3 reports the average predictive negative log-likelihood (NLL) for each method on the test data, which shows that DiGS significantly outperforms other baselines. We speculate that the performance gain comes from the fact that DiGS captures a broader range of modes.
@@
MALA $0.399\pm 0.014$ $5.0\times 10^{6}$
HMC $0.315\pm 0.012$ $5.0\times 10^{6}$
PT $0.241\pm 0.005$ $5.0\times 10^{6}$
DiGS $\mathbf{0.189\pm 0.002}$ $5.0\times 10^{6}$
@@ 4.3 Molecular Configuration Sampling
Each method is initialized at the minimum energy configuration as in Midgley et al. (2023) and generates $10^{6}$ samples of molecular configurations for evaluation. MALA runs 1,000 Langevin steps per sample. HMC runs 1,000 leapfrog steps per sample. DiGS uses the VP schedule with $T=2$ noise levels ($\alpha_{2}=0.1$ and $\alpha_{1}=0.9$), each with 100 Gibbs sweeps and 5 MALA denoising sampling steps per Gibbs sweep. Note that PT is the gold-standard method in molecular dynamics (MD) simulation, which serves as the ground-truth. The PT samples are taken from Midgley et al. (2023), which is generated using 21 chains starting at temperature 300K and increasing the temperature by 50K for each subsequent chain, where each chain is constructed by an HMC sampler with 1,000 leapfrog steps per sample. We follow Midgley et al. (2023) and treat $10^{7}$ PT samples as the ground-truth. In addition, we consider a baseline of $10^{6}$ PT samples as a reference. The quality of the sampled configurations is assessed by the Ramachandran plot (Ramachandran et al., 1963), which can be used to analyze how the protein folds locally. A Ramachandran plot is a 2D histogram for joint distribution of the two dihedral angles $\phi$ and $\psi$ in the bonds connecting an amino acid to the protein backbone, as shown in Figure 9. Table 4 shows the KL divergences for the Ramachandran plots $p(\phi,\psi)$ and the marginals $p(\phi)$ and $p(\psi)$ between the ground-truth samples and samples generated by each sampler. We can see that DiGS significantly outperforms MALA and HMC with the same number of energy evaluations. Moreover, DiGS also outperforms the gold-standard MD simulation method PT with 23×23$\times$23 × less energy evaluations. Figure 10 shows a visual comparison between the ground-truth Ramachandran plots and the one produced by DiGS, confirming that DiGS captures all modes with the correct weightings.
@@ 5 Conclusion
Diffusive Gibbs Sampling (DiGS) is an innovative sampler that combines recent development in diffusion models with a novel Metropolis-within-Gibbs scheme. DiGS offers significant improvements in sampling multi-modal distributions, surpassing traditional methods in both efficiency and accuracy. Its applicability across various fields (e.g., Bayesian inference and molecular dynamics) showcases the potential of DiGS to facilitate sampling complex distributions in numerous scientific applications.
5.1 Limitations and Future Work
[…]
Acknowledgements
We thank Andi Zhang for useful discussions on RDMC, Ruqi Zhang for highlighting Entropy-MCMC, an anonymous reviewer for discussions on PT, and Arnaud Doucet, Michael Hutchinson and Sam Power for pointing out proximal samplers.
MZ and DB acknowledge funding from the Cisco Centre of Excellence. WC acknowledges funding via a Cambridge Trust Scholarship (supported by the Cambridge Trust) and a Cambridge University Engineering Department Studentship (under grant G105682 NMZR/089 supported by Huawei R&D UK). JMHL acknowledges support from a Turing AI Fellowship under grant EP/V023756/1.
Impact Statement
The goal of this paper is to advance the field of machine learning. There could be many potential positive societal consequences of this methodological work, too numerous to specifically highlight here.
@@ Acknowledgements
$\log p_{\beta}(x=0)$ $=\beta\log p(x=0)+C_{\beta}$
$=\beta\left(-\frac{\mu^{2}}{2\sigma_{g}^{2}}-\log\sigma_{g}\right)+C_{\beta}.$ (21)
@@
Appendix C Discussions of Convergence Guarantees
In some specific cases, the convergence rate of DiGS may be established. Specifically, the convergence rate of DiGS depends on both the convergence of the inner denoising sampling step and the convergence of the outer Gibbs sampler. For instance, when using MALA to sample from the denoising posterior $p(x|\tilde{x})$, the convergence of MALA has been studied under different conditions such as strong log-concavity and log-Sobolev inequality (Vempala & Wibisono, 2019). Compared to direct sampling from the target distribution, the L2 regularizer in the log denoising posterior can improve the log-Sobolev condition:
$\log p(x|\tilde{x})=-E(x)-\frac{\lVert\alpha x-\tilde{x}\rVert^{2}}{2\sigma^{2}}+const,$ (23)
which in turn accelerates the convergence of MALA (Huang et al., 2023). Intuitively, this makes the target distribution more “Gaussian-like”. The convergence of the outer Gibbs sampler can be analyzed following the approach in Chen et al. (2022), demonstrating convergence under conditions of strong log-concavity/log-concavity under Wasserstein distance and log-Sobolev inequality under KL divergence. Therefore, the theoretical convergence guarantee of DiGS with a single noise level can be obtained by combining the convergence analyses of the inner MALA denoising sampling step and the outer Gibbs sampler from Vempala & Wibisono (2019); Chen et al. (2022); Huang et al. (2023).
However, in addition to the aforementioned factors, the multi-level noise schedule also plays a part in accelerating the convergence speed. Furthermore, our empirical investigation revealed that the performance of DiGS for practical problems such as molecular dynamics could also be affected by the initial condition (e.g., we followed the convention of molecular dynamics to initialize DiGS at the minimum energy configuration, which turns out to be better than random initialization). These analyses are out of the scope of this work, and we leave the theoretical study of the general DiGS framework as a future research direction.
Appendix D Experimental Details
In all experiments, the step sizes of MALA and HMC are tuned via trial-and-error so that the acceptance rates are close to optimal values $0.574$ (Roberts & Rosenthal, 1998) and $0.65$ (Neal et al., 2011), respectively.
D.1 Comparison of Initialization Strategies for the Denoising Sampling Step
D.2 Comparison of Gaussian Convolution Hyperparameters
D.3 Multi-Level Noise Scheduling
For the multi-level noise experiment with the VP schedule in Section 2.4, we define $\alpha_{t}=\alpha_{T}+(T-t)\Delta_{\alpha}$, where $\Delta_{\alpha}=(\alpha_{1}-\alpha_{T})/(T-1)$ and $\sigma_{t}=\sqrt{1-\alpha_{t}^{2}}$. We set $\alpha_{1}=0.9$ and $\alpha_{T}=0.1$, and experiment with $T\in\{2,3,4,5\}$ for the comparison shown in Figure 4(c). For evaluation, we employ the Maximum Mean Discrepancy (MMD) (Gretton et al., 2012) that utilizes 5 kernels with bandwidths $\{2^{-2},2^{-1},2^{0},2^{1},2^{2}\}$.
D.4 Comparison with Parallel Tempering
For comparison with parallel tempering on the “Mixture of Delta” problem in Figure 6 in Section 3.1, each sampler is initialized at the origin and generates 1,000 samples. DiGS employs a single noise level Gaussian convolution with parameters $\alpha=1$ and $\sigma=1$, complemented by 1,000 Gibbs sweeps per sample. The denoising sampling step in each Gibbs sweep begins with MH initialization, followed by 5 steps of MALA with a step size of $1{\times}10^{-3}$. PT consists of 5 chains with temperatures $\tau{=}\{1.0,5.62,31.62,177.83,1000.0\}$, where each chain is constructed by an HMC sampler with 1,000 leapfrog steps per sample and a step size of $1{\times}10^{-2}$.
[…]
D.5 Comparison with RDMC
For the comparison with Reverse Diffusion Monte Carlo (RDMC) in Figure 7 in Section 3.4, we follow the original setting discussed in Huang et al. (2023) and use ULA for generating samples for the posterior distribution. The score function of the reverse distribution using $K=5$ to construct the score approximation
$\nabla_{x_{t}}\log p(x_{t})\approx\left(\frac{\alpha}{K}\sum_{k=1}^{K}x_{0}^{(t,k)}-x_{t}\right)/\sigma_{t}^{2},$ (24)
where $x_{0}^{(t,k)}\sim p(x_{0}|x_{t})$ denotes the samples drawn using ULA with $L_{ULA}=5$ steps and a step size of $1{\times}10^{-2}$. These samples are initialized through importance sampling (IS) with $S_{is}=100$ samples. The discretization step size $\eta$ is set to $\Gamma/T$, with the scaling factor $a_{t}=e^{-(\Gamma-t\cdot\eta)}$ and the standard deviation $\sigma_{t}=\sqrt{1-\alpha_{t}^{2}}$. Following the procedure in Huang et al. (2023), we further apply ULA for additional $L_{ULA}$ steps after RDMC. The parameter $\Gamma$ is fixed at $0.1$ for our experiments, and we explore different values of $T\in\{1,2,3,4\}$, necessitating $TK(L_{ULA}+S_{is})+L_{ULA}$ energy evaluations to generate a single sample.
D.6 Mixture of 40 Gaussians
D.7 Bayesian Neural Networks
For the BNN experiment in Section 4.2, all methods are initialized at the same random sample from the prior and generate $150$ samples from the posterior $p(\theta|D_{\text{train}})$ for evaluation. MAP runs $7.5\times 10^{5}$ full-batch gradient descent steps with a step size of $3\times 10^{-2}$. MALA runs 5,000 Langevin steps per sample with a step size of $1\times 10^{-4}$. HMC runs 5,000 leapfrog steps per sample with a step size of $5\times 10^{-4}$. PT consists of 5 chains with temperatures $\tau=\{1.0,5.62,31.62,177.83,1000.0\}$, where each chain is constructed by an HMC sampler with 1,000 leapfrog steps per sample with a step size of $5\times 10^{-4}$. DiGS uses the VP schedule with $T=5$ noise levels, ranging from $\alpha_{T}=0.1$ to $\alpha_{1}=0.9$, each with 100 Gibbs sweeps and 10 MALA denoising sampling steps per Gibbs sweep with a step size of $1\times 10^{-4}$.
D.8 Molecular Configuration Sampling
For the molecular configuration sampling experiment in Section 4.3, each method is initialized at the minimum energy configuration as in Midgley et al. (2023) and generates $10^{6}$ samples of configurations for evaluation. MALA runs 1,000 Langevin steps per sample with a step size of $1\times 10^{-4}$. HMC runs 1,000 leapfrog steps per sample with a step size of $1\times 10^{-3}$. DiGS uses the VP schedule with $T=2$ noise levels ($\alpha_{2}=0.1$ and $\alpha_{1}=0.9$), each with 100 Gibbs sweeps and 5 MALA denoising sampling steps per Gibbs sweep with a step size of $1\times 10^{-4}$. Note that PT is the gold-standard method in molecular dynamics (MD) simulation, which serves as the ground-truth. The PT samples are taken from Midgley et al. (2023), which is generated using 21 chains starting at temperature 300K and increasing the temperature by 50K for each subsequent chain, where each chain is constructed by an HMC sampler with 1,000 leapfrog steps per sample. We follow Midgley et al. (2023) and treat $10^{7}$ PT samples as the ground-truth. In addition, we consider a baseline of $10^{6}$ PT samples as a reference.
Table 5: Sample quality on a 200-dimensional mixture of Gaussians problem.
Sampler MMD #energy evaluations
MALA $0.123\pm 0.035$ $10^{6}$
HMC $0.104\pm 0.025$ $10^{6}$
PT $0.094\pm 0.014$ $10^{6}$
DiGS $\mathbf{0.027\pm 0.010}$ $10^{6}$
Appendix E Additional Experimental Results
E.1 200-Dimensional Mixture of Gaussians
We take the MoG target in Figure 2(a) and generalize it to the 200-dimensional space. For each compared sampler, we draw 1,000 samples within a given compute budget of $10^{6}$ energy evaluations. All samplers are tuned to achieve their optimal acceptance rate as described in Section 4. Table 5 shows that DiGS significantly outperforms all the other samplers even in the high-dimensional space.
E.2 Molecular Configuration Sampling
The Ramachandran plots produced by all compared methods for the molecular configuration sampling experiment in Section 4.3 can be found in Figure 11.